# thanks: Corresponding author

### v1

@@ Abstract
In this paper, we systematically investigate the optical appearance of a Schwarzschild black hole illuminated by three geometrically thin accretion disk models under varying observational inclination angles. Based on the geometric relationship between the black hole and observer, we first divide the accretion disk into co-side and counter-side semi-disks. Subsequently, we analyze their differences in photon orbit distributions, transfer functions, and received intensities. Our results reveal distinct inclination-dependent behaviors between the two semi-disks: when the inclination approaches $\pi/2$, the co-side semi-disk develops richer ring structures with the lensed emission dominating the total brightness. Furthermore, through explicit specification of the emission profiles of the three models, we present optical images for both optically thin and thick disk scenarios at different inclinations. The results demonstrate that: (i) the bright rings in all three models become progressively compressed and deviate from circularity as inclination increases; (ii) for thick disks, partial rings are obscured and the overall intensity is lower than thin disks. These results may advance our understanding of general black hole imaging processes and provide a new approach to test gravitational theories through optical morphology studies.
@@ I Introduction
As one of the most influential theories of gravity, general relativity (GR) has profoundly influenced our understanding of the nature of gravity. In weak-field regions, the validity of GR has been confirmed by numerous classical experiments [1]. Moreover, in strong-field regions, the successful detection of gravitational waves (GWs) from a binary black hole (BH) merger [2] and the release of the first image of the supermassive BH [3] provided further confirmation of GR’s predictions. The released BH image of M87* reveals a distinct bright ring encircling a central dark region. This luminous structure arises from photons that are deflected by the BH’s extreme gravity and redirected toward the observer, whereas the dark shadow area corresponds to photons that have fallen into the BH. The critical photon orbit (photon sphere) forms the boundary between the bright ring and the dark shadow. The shadow image of a BH reflects the properties of its surrounding spacetime geometry. By analyzing the shadow, one can extract key parameters of the BH [4, 5], which has drawn significant attention to BH shadow research. Synge was the first to theoretically analyze the formation of shadows for Schwarzschild BH [6]. Later, Bardeen investigated the shadow of Kerr BH [7], finding that the Kerr BH’s shadow deviates from the circular shape of Schwarzschild’s case, with the distortion increasing as the spin parameter grows. Since then, extensive theoretical studies have been conducted on BH shadows in various gravitational theories and under different matter-field couplings [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
@@
Furthermore, considering astrophysical BHs in the universe, their surroundings inevitably contain various matter components. Under the influence of the BH’s strong gravitational field, this matter is accreted, forming different types of accretion structures around the BH [19]. Consequently, this leads to diverse BH shadow and bright ring configurations. Based on their geometric thickness and optical properties, accretion disks around BHs can be classified into thin and thick disks. The geometric thickness is typically quantified by the ratio of the disk thickness $D$ to its radius $R$ [20]. A disk is considered geometrically thin if $D/R\ll 1$. The optical thickness relates to the disk’s transparency. A disk is optically thick if it prevents the transmission of photons. Moreover, the distinct imaging features produced by different accretion disk models enable the verification of various gravitational theories through the optical appearance of BH accretion systems. In contrast to studies of geometrically thick disks [21], current research focus predominantly lies on geometrically thin disks. The first simulated image of a dynamically geometrically thin and optically thick accretion disk surrounding a Schwarzschild BH was produced by Luminet [22]. This pioneering work incorporated the observer’s inclination angle relative to the accretion disk, established angular relationships through spherical geometry, and ultimately generated a contour map on the observer’s plane displaying both the primary and secondary images of the BH. Subsequently, the study of BH imaging with such accretion disk models as emission sources, particularly under inclined viewing angles, has garnered widespread attention across various gravitational theories [20, 23, 24, 25, 26, 27, 28, 29, 30].
@@
In this study, we examine three geometrically thin accretion disk models under both optically thin and optically thick conditions, and investigate the optical appearance characteristics of Schwarzschild BH illuminated by these distinct types of accretion disks. Specifically, we divide the accretion disk into co-side and counter-side semi-disks based on their positions relative to the observer and BH. Using this dual-semi-disk framework, we analyze the photon orbit number distributions and light-ray intersection patterns for both semi-disks at various inclinations. Then we introduce corresponding transfer functions to establish the positional relationship between the BH and observer’s plane, which are then used to calculate the received intensities. Finally, we obtain optical images for both thin and thick disk scenarios under the three static accretion disk models.
The structure of this paper is organized as follows. In Section II, we briefly recall photon trajectories around a Schwarzschild BH. In Section III, based on the relative positions between the BH and observer, we divide the accretion disk into two semi-disks for separate analysis, examining photon trajectory intersections with both semi-disks and introducing their respective transfer functions. In Section IV, we obtain the optical appearance of a Schwarzschild BH illuminated by three types of optically thin accretion disk models at various observational inclination angles. Section V presents the corresponding optical images for the optically thick disk scenarios of all three models, with systematic comparisons to their thin-disk counterparts. Finally, conclusions are provided in Section VI. Throughout this paper, we adopt geometric units with $G=c=1$ and set the BH mass $M=1$ for computational purposes.
@@ II A Brief Recall of Photon Trajectories in Schwarzschild Black Hole Spacetime
${\rm d}s^{2}=-f(r){\rm d}t^{2}+\frac{1}{f(r)}{\rm d}r^{2}+r^{2}\left({\rm d}\theta^{2}+\sin^{2}\theta{\rm d}\phi^{2}\right),$ (1)
@@
Due to spherical symmetry, we restrict our analysis to photon trajectories confined to the equatorial plane ($\theta=\pi/2$). For photons, their geodesic tangent vectors $\left(\partial/\partial\tau\right)^{a}$ satisfy [57, 58]
@@
where $\tau$ is the affine parameter along the null geodesic. Within the Schwarzschild coordinate system, it follows that:
$\begin{split}0&=g_{ab}\left(\frac{\partial}{\partial\tau}\right)^{a}\left(\frac{\partial}{\partial\tau}\right)^{b}\\ &=-f(r)\left(\frac{{\rm d}t}{{\rm d}\tau}\right)^{2}+\frac{1}{f(r)}\left(\frac{{\rm d}r}{{\rm d}\tau}\right)^{2}+r^{2}\left(\frac{{\rm d}\phi}{{\rm d}\tau}\right)^{2}.\end{split}$ (4)
Note that in the Schwarzschild spacetime, there exist two Killing vector fields, $\left(\partial/\partial t\right)^{a}$ and $\left(\partial/\partial\phi\right)^{a}$, which correspond to the conserved quantities energy $E$ and angular momentum $L$, respectively. We have [57, 58]
$\begin{split}&E:=-g_{ab}\left(\frac{\partial}{\partial t}\right)^{a}\left(\frac{\partial}{\partial\tau}\right)^{b}=f(r)\frac{{\rm d}t}{{\rm d}\tau},\\ &L:=g_{ab}\left(\frac{\partial}{\partial\phi}\right)^{a}\left(\frac{\partial}{\partial\tau}\right)^{b}=r^{2}\frac{{\rm d}\phi}{{\rm d}\tau}.\end{split}$ (5)
@@
$\begin{split}\left(\frac{{\rm d}r}{{\rm d}\phi}\right)^{2}=r^{4}\left[\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}\right].\end{split}$ (6)
Here, $b\equiv L/E$ represents the photon’s impact parameter. For subsequent discussion, we reformulate the above expression using the substitution $u=1/r$, yielding
$\begin{split}\left(\frac{{\rm d}u}{{\rm d}\phi}\right)^{2}=\frac{1}{b^{2}}-f(\frac{1}{u})u^{2}.\end{split}$ (7)
Given a specific impact parameter $b$, its trajectory can be determined through Eq. (6). Figure  1 schematically displays the photon trajectories around the BH (represented by the black disk) for different impact parameters. Specifically, the red curve corresponds to photons with $b=b_{c}$ forming an unstable circular orbit around the BH due to gravitational deflection, where $b_{c}$ is referred to as the critical impact parameter. The black curves depict trajectories of photons with $b<b_{c}$ that ultimately fall into the event horizon, while the green curves represent photons with $b>b_{c}$ that are deflected by the BH and escape back to infinity.
@@
Notably, due to spherical symmetry, photon trajectories around a Schwarzschild BH can always be considered as confined to a plane passing through the BH. In such a plane, the aforementioned unstable circular orbits exist. These unstable circular orbits collectively form an unstable photon sphere around the BH, where perturbed photons either fall into the BH or escape to infinity. Consequently, without considering emission from accretion disks, observers would detect a perfectly circular BH shadow silhouette from any viewing angle. However, when considering luminous accretion disks, the optical appearance of the BH becomes dependent on both the disk type and observer’s inclination angle. We now examine the optical appearance of BHs surrounded by accretion disks at various inclination angles.
@@ III.1 Geometry of the black hole-observer system
For the Schwarzschild BH, we consider the presence of an optically thin and geometrically thin disk surrounding it. The disk emits radiation isotropically and is observed by distant observers at various inclination angles. The geometric relationship between the BH’s accretion disk and the observer is illustrated in Fig. 2.
Figure 2: The schematic diagram of the observer and BH coordinate systems, see [22, 29].
We assume the accretion disk lies strictly in the BH’s equatorial plane, with the observer positioned at infinity. The observer’s coordinate axis $O^{\prime}Y^{\prime}$ is coplanar with the BH’s axis $OM$, and the angle between $OZ$ and $OO^{\prime}$ defines the viewing inclination angle $\omega$. A photon emitted from point $P$ on the disk undergoes a deflection angle $\phi$ before reaching point $P^{\prime}$ on the observer’s plane. By projecting the observer’s plane $X^{\prime}O^{\prime}Y^{\prime}$ onto the BH coordinate system to obtain the $XOY^{\prime\prime}$ plane, we apply the spherical sine theorem to triangles $\triangle PXM^{\prime}$ and $\triangle PYM$ to establish the relationship between the deflection angle $\phi$, the viewing inclination $\omega$, and the polar angle $\eta$ in the observer’s coordinate system [22]:
@@
Through further derivation, we obtain
$\phi=\arctan{(\frac{1}{\tan{\omega}\sin{\eta}})}.$ (9)
To facilitate subsequent discussion, we partition the BH’s accretion disk along the $XOY^{\prime\prime}$ plane into two distinct regions: the co-side semi-disk (located on the same side as the observer, to the right of the $XOY^{\prime\prime}$ plane) and the counter-side semi-disk (on the opposite side, to the left of the $XOY^{\prime\prime}$ plane).
@@ III.2 Classification of photon trajectories
Photons emitted from the accretion disk and reaching the observer can equivalently be treated as backward-traced light rays from the observer intersecting with the disk near the BH. Each intersection corresponds to brightness contribution from the disk. Since light rays may intersect the accretion disk multiple times, we follow the conventional classification scheme of photon trajectories from [31], defining direct emission for rays intersecting the disk once, lensed emission for twice, and photon ring emission for three or more times by computing the orbit number $n=\phi/{(2\pi)}$, where $\phi$ represents the light deflection angle. Combining with Eq. (9), we derive:
$n=\frac{\phi}{2\pi}=\frac{1}{{2\pi}}\arctan{(\frac{1}{\tan{\omega}\sin{\eta}})}.$ (10)
@@
Subsequently in Fig. 3, we demonstrate the intersection patterns between light rays (emitted from a distant observer at infinity on the right side) and the accretion disk around a Schwarzschild BH at non-zero inclination angles, specifically showing the counter-side semi-disk (left panel) and co-side semi-disk (right panel). The green dashed line $NY^{\prime\prime}$ represents the $XOY^{\prime\prime}$ plane from Fig. 2, while the blue solid line indicates the equatorial plane of the accretion disk. Black, orange, and red curves correspond to light rays intersecting the disk once, twice, and three or more times, respectively. We observe distinct intersection patterns between the counter-side semi-disk and co-side semi-disk, which motivates separate discussions of the orbit numbers $n_{1}$ for the counter-side semi-disk and $n_{2}$ for the co-side semi-disk.
@@
Figure 3: Light trajectories from a distant right-side observer with inclination, intersecting the counter-side semi-disk (left) and co-side semi-disk (right) of a Schwarzschild BH. The blue solid curve represents the accretion disk, while the black, orange, and red curves correspond to light trajectories intersecting the disk once, twice, and three or more times, respectively.
@@
In Figs. 4 and 5, we present the variations of orbit numbers $n_{1}$ and $n_{2}$ with impact parameter $b$ for different values of $\omega$ and $\eta$, respectively. In Fig. 4, with the polar angle fixed at $\eta=\pi/2$, the impact parameter range corresponding to the lensed emission region of the counter-side semi-disk gradually decreases with increasing viewing inclination, while that of the co-side semi-disk shows significant enlargement. Figure 5 demonstrates similar trends at fixed inclination $\omega=17\pi/36$ across various polar angles $\eta$. Additionally, we find the orbit numbers remain invariant under inclination changes when $\eta=0$.
@@
We emphasize that for any given observer polar angle $\eta$, light rays intersect both semi-disks simultaneously as shown in Fig. 3. Moreover, the observed image exhibits exact symmetry about the $O^{\prime}Y^{\prime}$ axis. This symmetry reduces our analysis to the range $20\leq\eta\leq\pi/$ 2 while fully characterizing the entire image. All subsequent $\eta$ values are implicitly constrained to this angular domain without repeated specification.

### v2

@@ Abstract
In this paper, we systematically investigate the optical appearance of a Schwarzschild black hole illuminated by three geometrically thin accretion disk models under varying observational inclination angles. Based on the geometric relationship between the black hole and observer, we first divide the accretion disk into co-side and counter-side semi-disks. We then analyze light ray trajectories, and calculate the total number of orbits and transfer functions for both semi-disks. The results reveal distinct inclination-dependence of lensed regions on different semi-disks: as inclination increases, the lensed region contracts for the counter-side semi-disk while expanding for the co-side one. Furthermore, through explicit specification of the emission profiles of the three models, we present optical images for both optically thin and thick disk scenarios at different inclinations. The results demonstrate that: (i) the bright rings in all three models become progressively compressed and deviate from circularity as inclination increases; (ii) for thick disks, partial rings are obscured and the overall intensity is lower than thin disks. These results may advance our understanding of general black hole imaging processes and provide a new approach to test gravitational theories through optical morphology studies.
@@ I Introduction
As one of the most influential theories of gravity, general relativity (GR) has profoundly influenced our understanding of the nature of gravity. In weak-field regions, the validity of GR has been confirmed by numerous classical experiments [1]. Moreover, in strong-field regions, the successful detection of gravitational waves (GWs) from a binary black hole (BH) merger [2] and the release of the first image of the supermassive BH [3] provided further confirmation of GR’s predictions. The released BH image of M87* reveals a distinct bright ring encircling a central dark region. This luminous structure arises from photons that are deflected by the BH’s extreme gravity and redirected toward the observer, whereas the dark shadow area corresponds to photons that have fallen into the BH. The shadow image of a BH reflects the properties of its surrounding spacetime geometry. By analyzing the shadow, one can extract key parameters of the BH [4, 5], which has drawn significant attention to BH shadow research. Synge was the first to theoretically analyze the formation of shadows for Schwarzschild BH [6]. Later, Bardeen investigated the shadow of Kerr BH, and found that the Kerr BH’s shadow deviates from the circular shape of Schwarzschild’s case, with the distortion increasing as the spin parameter grows [7]. Since then, extensive theoretical studies have been conducted on BH shadows in various gravitational theories, including matter couplings [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18].
@@
Furthermore, considering astrophysical BHs in the Universe, their surroundings inevitably contain various matter components. Under the influence of the BH’s strong gravitational field, this matter is accreted by the BH, forming different types of accretion structures around the BH [19]. Based on their geometric thickness and optical properties, accretion disks around BHs can be classified into thin and thick disks. The geometric thickness is typically quantified by the ratio of the disk thickness $D$ to its radius $R$ [19, 20], where a disk is considered geometrically thin if $D/R\ll 1$. In contrast, the optical thickness relates to the disk’s transparency; specifically, a disk is optically thick if it prevents the transmission of photons. The distinct imaging features produced by these different accretion disk models make it possible to test various gravitational theories through the optical appearances of BH accretion systems. Previous studies have extensively investigated both geometrically thick and thin disks [19, 21]. The first simulated image of a dynamically geometrically thin and optically thick accretion disk surrounding a Schwarzschild BH was produced by Luminet [22]. This pioneering work incorporated the observer’s inclination angle relative to the accretion disk, established angular relationships through spherical geometry, and ultimately generated a contour map on the observer’s plane displaying both the primary and secondary images of the BH. Subsequently, the study of BH imaging with such accretion disk models as emission sources, particularly under inclined viewing angles, has garnered widespread attention across various gravitational theories [20, 23, 24, 25, 26, 27, 28, 29, 30].
@@
In this study, we examine three geometrically thin accretion disk models under both optically thin and optically thick conditions, and investigate the optical appearance characteristics of a Schwarzschild BH illuminated by these distinct types of accretion disks. Specifically, we divide the accretion disk into co-side and counter-side semi-disks based on their positions relative to the observer and BH. Based on the structure of two semi-disks, we analyze the photon orbit number distributions and light-ray intersection patterns for both semi-disks at various inclinations. Then we introduce corresponding transfer functions to establish the positional relationship between the accretion disk of the BH and observer’s plane, which are then used to calculate the received intensities. Finally, we obtain optical images for both optically thin and thick disk scenarios under the three static accretion disk models.
The structure of this paper is organized as follows. In Sec. II, we briefly recall photon trajectories around a Schwarzschild BH. In Sec. III, based on the relative positions between the BH and observer, we divide the accretion disk into two semi-disks for separate analysis, examining photon trajectory intersections with both semi-disks and introducing their respective transfer functions. In Sec. IV, we obtain the optical appearance of a Schwarzschild BH illuminated by three types of optically thin accretion disk models at various observational inclination angles. Sec. V presents the corresponding optical images for the optically thick disk scenarios of all three models, with systematic comparisons to their thin-disk counterparts. Finally, a summary is provided in Sec. VI. Throughout this paper, we adopt geometric units with $G=c=1$ and set the BH mass $M=1$ for computational purposes.
@@ II A Brief Recall of Photon Trajectories in Schwarzschild Black Hole Spacetime
${\rm d}s^{2}=-f(r){\rm d}t^{2}+\frac{1}{f(r)}{\rm d}r^{2}+r^{2}\left({\rm d}\theta^{2}+\sin^{2}\theta{\rm d}\varphi^{2}\right),$ (1)
@@
Here $M$ denotes the mass of the BH. Due to spherical symmetry of spacetime, we restrict our analysis to photon trajectories confined to the equatorial plane ($\theta=\pi/2$). The tangent vector field $\left(\partial/\partial\tau\right)^{a}$ of the null geodesic satisfy [57, 58]
@@
where $\tau$ is the affine parameter along the null geodesic. In the Schwarzschild coordinate system, Eq. (3) can be written as
$\begin{split}0&=g_{ab}\left(\frac{\partial}{\partial\tau}\right)^{a}\left(\frac{\partial}{\partial\tau}\right)^{b}\\ &=-f(r)\left(\frac{{\rm d}t}{{\rm d}\tau}\right)^{2}+\frac{1}{f(r)}\left(\frac{{\rm d}r}{{\rm d}\tau}\right)^{2}+r^{2}\left(\frac{{\rm d}\varphi}{{\rm d}\tau}\right)^{2}.\end{split}$ (4)
Note that in the Schwarzschild spacetime, there exist two Killing vector fields, $\left(\partial/\partial t\right)^{a}$ and $\left(\partial/\partial\varphi\right)^{a}$, which determine two conserved quantities, the energy $E$ and angular momentum $L$, by [57, 58]
$\begin{split}&E:=-g_{ab}\left(\frac{\partial}{\partial t}\right)^{a}\left(\frac{\partial}{\partial\tau}\right)^{b}=f(r)\frac{{\rm d}t}{{\rm d}\tau},\\ &L:=g_{ab}\left(\frac{\partial}{\partial\varphi}\right)^{a}\left(\frac{\partial}{\partial\tau}\right)^{b}=r^{2}\frac{{\rm d}\varphi}{{\rm d}\tau}.\end{split}$ (5)
@@
$\begin{split}\left(\frac{{\rm d}r}{{\rm d}\varphi}\right)^{2}=r^{4}\left[\frac{1}{b^{2}}-\frac{f(r)}{r^{2}}\right].\end{split}$ (6)
Here, $b\equiv L/E$ represents the photon’s impact parameter. For subsequent discussion, we reformulate the above expression using the substitution $u=1/r$, and get
$\begin{split}\left(\frac{{\rm d}u}{{\rm d}\varphi}\right)^{2}=\frac{1}{b^{2}}-f\left(\frac{1}{u}\right)u^{2}.\end{split}$ (7)
Hence, given a specific impact parameter $b$ associated to a photon, the trajectory of the photon can be determined through Eq. (6) [or Eq. (7)]. Figure 1 schematically displays the photon trajectories around the BH (represented by the black disk) for different impact parameters. Specifically, the red curve corresponds to photons with $b=b_{c}$ forming an unstable circular orbit around the BH due to gravitational deflection, where $b_{c}$ is referred to as the critical impact parameter. The black curves depict trajectories of photons with $b<b_{c}$ that ultimately fall into the event horizon, while the green curves represent photons with $b>b_{c}$ that are deflected by the BH and escape back to infinity.
@@
Notably, due to spherical symmetry, it turns out that any one photon trajectory around a Schwarzschild BH can always be confined to a plane passing through the center of the BH. In such a plane, the aforementioned unstable circular orbits exist. These unstable circular orbits collectively form an unstable photon sphere around the BH, where perturbed photons either fall into the BH or escape to infinity. Hence, when a BH is illuminated from behind by a large, distant planar screen that emits isotropiclly with uniform brightness, a perfectly circular BH shadow silhouette is formed in the observer’s view at all inclination angles [31]. However, when considering luminous accretion disks, the optical appearance of the BH becomes dependent on both the disk type and observer’s inclination angle. In what follows, we will examine the optical appearance of BHs surrounded by accretion disks at various inclination angles.
@@ III.1 Geometry of the black hole-observer system
For the Schwarzschild BH, we consider the presence of an optically and geometrically thin disk as the light source surrounding it. The disk emits radiation isotropically and is observed by distant observers at various inclination angles. To study the appearance of a Schwarzschild BH surrounded by the disk, it is convenient to introduce a (local) Cartesian coordinate system $\{X,Y,Z\}$ based on the center of the BH in Fig. 2 [22, 29, 59]. Assuming that the BH center and a distant observer are located at point $O$ and $O^{\prime}$ respectively, and the accretion disk lies strictly in the equatorial plane ($\theta=\pi/2$) of the BH. We specify the disk’s symmetry axis as the $OZ$-axis, and the intersection between the equatorial plane (the blue plane in Fig. 2) and a plane passing through the $OZ$-axis and the radial line $OO^{\prime}$ as the $OY$-axis. The plane through the $OZ$-axis and $OO^{\prime}$ can be obtained by rotating the equatorial plane about $O$. Thus, the $OX$-axis in the equatorial plane is determined by requiring that the coordinate system $\{X,Y,Z\}$ is a right-handed one, in the other words, the $OX$-axis is obtained by rotating counterclockwise about the $OZ$-axis through $\pi/2$. Once the coordinate system $\{X,Y,Z\}$ is specified, the spherical coordinate system $\{r,\theta,\varphi\}$ adapted to the coordinate system $\{X,Y,Z\}$ is determined in the conventional sense. Hence, any one point in space can then be described by spherical coordinates $(r,\theta,\varphi)$, for instance, $O^{\prime}$ has coordinates $(r_{o},\omega,\pi/2)$ with $r_{o}\gg M$. The angle $\omega$ is called the inclination angle. Next, we introduce the observer’s plane which passes through the observer’s position $O^{\prime}$ and orthogonal (measured by the Schwarzschild metric) to the radial line $OO^{\prime}$. Furthermore, in this plane, we construct a coordinate system $\{X^{\prime},Y^{\prime}\}$ based on $O^{\prime}$, called as the observer’s coordinate system. The $O^{\prime}Y^{\prime}$-axis is specified as the intersection of the observer’s plane with the plane formed by the $OZ$-axis and the radial line $OO^{\prime}$, oriented in the direction of decreasing $\theta$. The $O^{\prime}X^{\prime}$-axis is aligned with the intersection of the observer’s plane and the plane through the $OX$-axis and the radial line $OO^{\prime}$, pointing in the direction of increasing $\varphi$. Moreover, we perform a counterclockwise rotation of the $\{X,Y,Z\}$ frame by $\pi/2-\omega$ about the $OX$-axis, which brings the $OZ$-axis into alignment with $OY^{\prime\prime}$. Thus, the plane $XOY^{\prime\prime}$ maintains a $\pi/2$ angle relative to $OO^{\prime}$.
Figure 2: The schematic diagram of the observer and BH coordinate systems. The BH center is located at point $O$, and the observer is positioned at point $O^{\prime}$. The viewing inclination is denoted by $\omega$. A photon emitted from point $P$ on the accretion disk arrives at point $P^{\prime}$ on the observer’s plane, undergoing a total azimuthal angle change $\angle POM\equiv\phi$, with $\eta$ representing the polar angle in the observer’s plane.
Consider a point $P$ with coordinates $(r_{P},\pi/2,\varphi_{P})$ in the accretion disk plane and a spherical surface with a radius $r_{P}$ centered at $O$. Denote the intersections of the surface with the coordinate axes $OX$ and $OY$ by points $x$ and $y$, respectively. The trajectory of a photon emitted from $P$ is confined to a plane through $O$ (the pink plane in Fig. 2), referred to as a trajectory plane. The intersections of these planes with the spherical surface are marked in Fig. 2. The photon ultimately reaches a point $P^{\prime}$ on the observer’s plane, forming a polar angle $\eta$ with the $O^{\prime}X^{\prime}$-axis while accumulating a total azimuthal angle change $\phi$ during propagation. By applying the spherical sine theorem to the spherical triangles, one can derive the fundamental relation among $\phi$, $\omega$, and $\eta$ as [22]
@@
Hence, we obtain
$\phi=\arctan\left(\frac{1}{\tan{\omega}\sin{\eta}}\right).$ (9)
To facilitate subsequent discussion, it is convenient to partition the BH’s accretion disk along $OX$-axis by the $XOY^{\prime\prime}$ plane (the green plane in Fig. 2) into two distinct semi-disks: the co-side semi-disk (located on the same side as the observer, to the right of the $XOY^{\prime\prime}$ plane) and the counter-side semi-disk (on the opposite side, to the left of the $XOY^{\prime\prime}$ plane).
@@ III.2 Classification of photon trajectories
Photons emitted from the accretion disk and reaching the observer can equivalently be treated as backward-traced light rays from the observer intersecting with the disk near the BH. Each intersection corresponds to brightness contribution from the disk. Since light rays may intersect the accretion disk multiple times in the case of optically thin, we follow the conventional classification scheme of photon trajectories proposed in [31], and define direct emission for rays intersecting the disk once, lensed emission for twice, and photon ring emission for three or more times by considering the orbit number $n=\phi/{(2\pi)}$, where $\phi$ represents the total change in orbital plane azimuthal angle outside the horizon. Combining with Eq. (9), we have
$n=\frac{\phi}{2\pi}=\frac{1}{{2\pi}}\arctan\left(\frac{1}{\tan{\omega}\sin{\eta}}\right).$ (10)
When photons arrive at the plane $XOY^{\prime\prime}$ from a distant observer, they undergo an azimuthal angle change of $\pi/2$. In Fig. 3, we single out the trajectory plane (the pink plane) in Fig. 2 and take the case where photons intersect each semi-disk once as an example, in order to show the dependence of the total changes $\phi$ on these two semi-disks. The two points $P_{2}$ and $P_{1}$ represent the intersection points of photons with the counter-side and co-side semi-disks, respectively. It is easy to see that from Fig. 3 one only needs to calculate the angle $\lambda$ between the plane $XOY^{\prime\prime}$ and the accretion disk within this trajectory plane to determine the total azimuthal angle of photons reaching the two semi-disks, and thus obtain the total change $\phi=\pi/2\pm\lambda$, where $+$ for the counter-side semi-disk while -- for the co-side semi-disk. Based on the geometric relations in Fig. 2, we get
$\lambda=\arctan{(\tan{\omega}\sin{\eta})}.$ (11)
Figure 3: Schematic of geometric relations in the photon trajectory plane, where $O$ is the BH position and $O^{\prime}$ corresponds to the center of the observer’s plane. The blue dashed line represents the accretion disk, while the green dashed line indicates the plane $XOY^{\prime\prime}$.
@@
In Fig. 4, we demonstrate the intersection patterns between light rays (emitted from a distant observer at infinity on the right side) and the accretion disk around a Schwarzschild BH at non-zero inclination angles, specifically showing the counter-side semi-disk (left panel) and co-side semi-disk (right panel). In Fig. 4, the green dashed line $NY^{\prime\prime}$ represents the $XOY^{\prime\prime}$ plane in Fig. 2, while the blue dashed line indicates the equatorial plane of the accretion disk. Black, orange, and red curves in Fig. 4 denote direct, lensed, and photon ring trajectories, respectively. The distinct intersection patterns between the counter-side semi-disk and co-side semi-disk motivates us to take separate treatments of the orbit numbers for the two semi-disks. In what follows, we denote the orbit numbers for the counter-side and the co-side semi-disks by $n_{1}$ and $n_{2}$, respectively.
@@
Figure 4: Light trajectories from a distant right-side observer with inclination $\omega$, intersecting the counter-side semi-disk (left) and co-side semi-disk (right) of a Schwarzschild BH. The black, orange, and red curves correspond to direct, lensed, and photon ring trajectories, respectively. The blue dashed curve depicts the accretion disk, while the green dashed line represents the $XOY^{\prime\prime}$ plane.
@@
In Figs. 5 and 6, we present the variations of orbit numbers $n_{1}$ and $n_{2}$ with impact parameter $b$ for different values of $\omega$ and $\eta$, respectively. In Fig. 5, with the polar angle fixed at $\eta=\pi/2$, the impact parameter range corresponding to the lensed emission region of the counter-side semi-disk gradually decreases with increasing viewing inclination, while that of the co-side semi-disk shows significant enlargement. Figure 6 demonstrates similar trends at fixed inclination $\omega=17\pi/36$ across various polar angles $\eta$. Additionally, we find the orbit numbers remain invariant under inclination changes when $\eta=0$.
@@
It is crucial to emphasize that for a given polar angle $\eta$ in the observer’s plane, the counter-side and co-side semi-disks are imaged in the upper and lower half-planes of the $O^{\prime}X^{\prime}$ axis in the observer’s coordinate system, respectively, as shown in Fig. 4. Moreover, the observed image exhibits exact symmetry about the $O^{\prime}Y^{\prime}$ axis. This symmetry reduces our analysis to the range $20\leq\eta\leq\pi/$ 2 while fully characterizing the entire image. All subsequent $\eta$ values are implicitly constrained to this angular domain without repeated specification.